OpenStudy (anonymous):

Ok for example if i have the problem square root(180) plus square root(180) Do i add the 180 together to 360. Or do I simplify first and then add the fives that results in the radical. Or do I just leave 180 as it is which would then make it 2 squareroot(180)

OpenStudy (anonymous):

\[\sqrt{180}+\sqrt{180}\]

OpenStudy (callisto):

\[\sqrt{180}+\sqrt{180} \ne \sqrt{360}\]Just let \(a=\sqrt{180}\), then you get a+a for your question, which becomes 2a. So, you get \(2\sqrt{180}\). But that's not the end of the story. You can still simplify \(\sqrt{180}\) as \(\sqrt{180}=\sqrt{9\times 4 \times 5}=\sqrt{6^2\times5}\)

OpenStudy (anonymous):

OK thank you that helps a whole lot i just forgot what i had to do. Your awesome!

OpenStudy (callisto):

You're welcome :)